Jefferson Lab HALL A Logbook for September 05 - All EXCEPT Automatic Entries

  • Main INDEX

    Wednesday

  • 156342 09/28/05 15:20 H. Benaoum Beam Polarization Measured by Compton Polarimeter after Labor day Run
  • 156341 09/28/05 12:00 H. Benaoum Raw, Regressed, Dithering Corrected and Regression Corrected Detectors Normalized Asymmetries versus Slug Number (1-121) without slugs 39,40 and 41 .
  • 156340 09/28/05 10:16 H. Benaoum WAC Daily Summary for Sept. 26-28
  • 156339 09/28/05 09:12 H. Benaoum Raw and Regressed Detectors Normalized Asymmetries versus Slug Number (77-121 ) after transverse polarization flip.
  • 156338 09/28/05 09:06 H. Benaoum WAC Summary for Slug 121.
  • 156337 09/28/05 08:03 H. Benaoum WAC Summary for Slug 120
  • 156336 09/28/05 07:11 R. Michaels bmw and feedback off
  • 156334 09/28/05 07:07 R. Michaels owl summary, 4He ended
  • 156330 09/28/05 06:17 R. Michaels spot++
  • 156329 09/28/05 06:16 R. Michaels target screens
  • 156326 09/28/05 05:21 R. Michaels Q2 scan results
  • 156325 09/28/05 05:08 R. Michaels magnet screen, run 5439
  • 156323 09/28/05 05:03 R. Michaels RQ2 restored, will start production
  • 156322 09/28/05 04:41 R. Michaels flipper. wanted prod, but RQ2 trip.
  • 156321 09/28/05 04:30 R. Michaels Q2 scan, part III
  • 156318 09/28/05 04:15 lkaufman Scanner Contour plots
  • 156308 09/28/05 03:01 R. Michaels Q2 scan studies part II
  • 156299 09/28/05 01:53 R. Michaels more Q2 scanner studies, part I
  • 156292 09/28/05 01:13 R. Michaels production run 5424
  • 156287 09/28/05 00:48 R. Michaels correction to swing report
  • 156286 09/28/05 00:28 R. Michaels system recovery, tonights plan
  • 156285 09/28/05 00:25 G.M. Urciuol swing shift report
  • Tuesday

  • 156281 09/27/05 23:15 bertin lost control all the magnets right,left,septa...
  • 156279 09/27/05 22:57 G.M. Urciuol HallaATools Screen
  • 156278 09/27/05 22:50 pking Adjusting JT valve
  • 156276 09/27/05 22:33 Adaq spot++ check on LHe tgt
  • 156273 09/27/05 22:26 pking Moving the He target back; opened JT valve to regular settings
  • 156272 09/27/05 22:19 Moffit Late Entry: Profile scintillators back to HRS DAQ
  • 156271 09/27/05 22:18 G.M. Urciuol Profile scan with Carbon target
  • 156269 09/27/05 21:25 Adaq spot++ check on carbon target
  • 156265 09/27/05 21:17 pking Moved to the C target
  • 156264 09/27/05 21:12 G.M. Urciuol Profile scan with BeO target
  • 156263 09/27/05 21:12 pking Closed JT valve during BeO running.
  • 156261 09/27/05 20:23 pking Move target to BeO at 19:50
  • 156260 09/27/05 20:22 pking Target HPH dropped to zero; hall access to replace HPH power supply.
  • 156257 09/27/05 20:03 pking spot++ check on BeO
  • 156252 09/27/05 17:02 H. Benaoum Structure of Profile Scans with 20 runs Combined
  • 156247 09/27/05 16:38 pking Compton run 12312
  • 156246 09/27/05 16:16 pking Changed the LOW alarm limit for the HPH
  • 156245 09/27/05 16:02 H. Benaoum Helium Profile Scans with 4 Densities
  • 156242 09/27/05 15:29 S. Fuchs Hall A Tools screen
  • 156241 09/27/05 15:26 S. Fuchs target screen
  • 156223 09/27/05 10:33 E.Chudakov Hall C bleedthrough 3
  • 156222 09/27/05 10:03 E.Chudakov Hall C bleedthrough 2
  • 156221 09/27/05 10:01 E.Chudakov Hall C bleedthrough 1
  • 156216 09/27/05 08:37 R. Michaels scanner was left in (affects 5409)
  • 156214 09/27/05 08:12 R. Michaels run 5409 ended
  • 156213 09/27/05 08:10 R. Michaels Right septum quench
  • 156210 09/27/05 08:01 R. Michaels owl summary
  • 156209 09/27/05 07:57 bertin hp power low
  • 156208 09/27/05 07:56 bertin target screen
  • 156207 09/27/05 07:52 Adaq tools
  • 156205 09/27/05 07:31 R. Michaels spot++
  • 156202 09/27/05 07:27 R. Michaels Q2 scanner alignment
  • 156184 09/27/05 06:29 R. Michaels Q2 versus rate
  • 156175 09/27/05 05:28 R. Michaels return to nominal P_L
  • 156173 09/27/05 05:27 R. Michaels dipole scan
  • 156169 09/27/05 05:09 R. Michaels run 2776 was 12%
  • 156160 09/27/05 03:56 R. Michaels Q2 results
  • 156157 09/27/05 03:19 R. Michaels HV scan
  • 156139 09/27/05 02:19 Moffit Slices in focal plane, HRS run 2689
  • 156131 09/27/05 01:59 R. Michaels Spectr. run 2760. Tonight's run plan.
  • 156128 09/27/05 01:35 Armstrong target returned to 100% density
  • 156127 09/27/05 01:31 R. Michaels target refill, scanner scint to spectr. DAQ
  • 156126 09/27/05 01:27 lkaufman Detector/BCM Linearity
  • 156125 09/27/05 00:17 lkaufman 2D plots for profile scans
  • 156124 09/27/05 00:00 lkaufman Scan comparisons for low density 4He runs
  • Monday

  • 156123 09/26/05 23:56 pking Reset the target density parameter in the PID loop
  • 156122 09/26/05 23:51 pking Target vacuum has come back down
  • 156121 09/26/05 23:48 pking Shift summary for the first part of swing shift (16:00-18:00)
  • 156120 09/26/05 23:45 Nilanga Shift Summary part II
  • 156118 09/26/05 23:33 Yi Qiang Hall A Tool Screen
  • 156116 09/26/05 23:16 Yi Qiang Profile Scan 5408
  • 156112 09/26/05 23:00 pking Target screens with 38 uA on the 49 psi target
  • 156110 09/26/05 22:50 Armstrong reduced target density to around 40% of nominal
  • 156109 09/26/05 22:50 Armstrong reduced target density to around 40% of nominal
  • 156108 09/26/05 22:49 pking Changed target density parameter; target charts at 48 psi
  • 156107 09/26/05 22:23 pking Hall access to reduce the target pressure to 45 psi.
  • 156105 09/26/05 22:12 pking Target vacuum spikes
  • 156101 09/26/05 21:39 pking Target vacuum has increased to 2.1e-6
  • 156100 09/26/05 21:38 Yi Qiang Profile Scan 5406
  • 156098 09/26/05 21:26 Yi Qiang Profile Scan 5405
  • 156095 09/26/05 21:06 Yi Qiang Profile scan 5404
  • 156089 09/26/05 20:46 pking Target loop parameters changed for the lower density target
  • 156086 09/26/05 20:35 pking Spot++ for the 78 psi target runs
  • 156083 09/26/05 20:26 Yi Qiang Profile Scan Plots
  • 156082 09/26/05 20:22 Armstrong reduced 4He density to 60%
  • 156081 09/26/05 20:12 pking Hall access to reduce the target pressure.
  • 156066 09/26/05 19:10 Nilanga VDC run 2750 :trigger rates
  • 156065 09/26/05 18:59 Yi Qiang Profile Scan Plots
  • 156060 09/26/05 18:31 Nilanga Run 5397 junk
  • 156055 09/26/05 18:11 pking Spot++ capture
  • 156052 09/26/05 18:02 stevensm techs/right septum quench on ramp
  • 156051 09/26/05 17:56 pking Right septum work is complete; ask for beam permit.
  • 156050 09/26/05 17:51 R. Feuerbach Estimate of HAP-de thresholds
  • 156047 09/26/05 17:10 pking Status of work on right septum magnet
  • 156046 09/26/05 16:57 pking Target screens at 106 psi
  • 156043 09/26/05 16:48 lkaufman Tungsten preradiator back in quartz scanner
  • 156042 09/26/05 16:43 Armstrong Reduced Helium target density by 23%
  • 156041 09/26/05 16:19 pking D. Armstrong enters the hall for target density reduction.
  • 156039 09/26/05 16:06 H. Benaoum Day Shift Summary
  • 156037 09/26/05 14:39 H. Benaoum WAC Summary for Slug 119
  • 156036 09/26/05 14:36 E.Chudakov Moller results
  • 156035 09/26/05 13:58 R. Snyder 4A & 4B position differences through 118 slugs
  • 156034 09/26/05 13:53 H. Benaoum Regression & Dithering Slopes for Detectors vs. Slug (1-118) without slugs 1,2,3,39,40 & 41.
  • 156033 09/26/05 13:39 lkaufman Scanner comparisons - w/ and w/o preradiator
  • 156032 09/26/05 13:36 H. Benaoum Right Septum quenched
  • 156031 09/26/05 13:12 lkaufman WAC Daily Summary for Sept. 25
  • 156030 09/26/05 13:05 lkaufman WAC Daily Summary for Sept. 23-24
  • 156029 09/26/05 11:21 H. Benaoum Raw, Regressed, Dithering Corrected and Regression Corrected Detectors Normalized Asymmetries versus Slug Number (1-118) without slugs 39,40 and 41
  • 156028 09/26/05 11:09 H. Benaoum Raw and Regressed Detectors Normalized Asymmetries versus Slug Number (77-118 ) after transverse polarization flip.
  • 156027 09/26/05 10:35 E.Chudakov Start Moller measurement
  • 156026 09/26/05 10:33 H. Benaoum Beam Polarization Measured by Compton Polarimeter from 12 till 26 September
  • 156025 09/26/05 10:20 H. Benaoum 86.5% Beam Polarization Measured by Compton Polarimeter ( run 12308
  • 156023 09/26/05 09:51 H. Benaoum Compton Quarter Wave Plate Scan
  • 156018 09/26/05 09:05 H. Benaoum Left Dipole Ramping
  • 156017 09/26/05 09:00 fansler techs
  • 156014 09/26/05 08:26 H. Benaoum Access to Hall to Reset Left Dipole
  • 156012 09/26/05 08:17 H. Benaoum Lost Left Dipole
  • 156009 09/26/05 08:03 Saha Shift Summary
  • 156008 09/26/05 07:26 kfuoti IHWP out
  • 156005 09/26/05 06:37 lkaufman WAC Daily Summary Sept. 23-24
  • 156000 09/26/05 05:29 lkaufman WAC Summary for Slug 118
  • 155999 09/26/05 05:28 lkaufman WAC Summary for Slug 117
  • 155998 09/26/05 05:26 lkaufman WAC Summary for Slug 116
  • 155997 09/26/05 05:22 lkaufman WAC Summary for Slug 115
  • 155996 09/26/05 05:09 lkaufman WAC Summary for Slug 114
  • 155995 09/26/05 05:07 lkaufman WAC Summary for Slug 113
  • 155994 09/26/05 04:50 kfuoti Compton Cavity Power
  • 155989 09/26/05 02:44 Fuoti/Kaufma 5384: Fine profile scan with scintillator and quartz
  • 155986 09/26/05 02:23 Saha Q2 Scan Run 5384
  • 155982 09/26/05 02:14 Saha Q2 Scan Run 5383
  • 155980 09/26/05 02:06 Saha Q2 Scan Run 5382
  • 155974 09/26/05 01:59 Moffit spot++ captured!
  • 155973 09/26/05 01:56 Saha Scanner HVs
  • 155970 09/26/05 01:39 Moffit Access activity
  • 155967 09/26/05 01:02 pking Target temperature camera has gone blank
  • 155966 09/26/05 00:30 pking Swing shift summary
  • 155965 09/26/05 00:24 pking Compton run 12303
  • 155961 09/26/05 00:00 pking Opening the spectrometer huts for profile scanner work
  • Sunday

  • 155960 09/25/05 23:47 pking Left dipole is at full field
  • 155959 09/25/05 23:39 pking Rest of Left dipole power supply
  • 155958 09/25/05 23:15 pking BRyan goes into Hall A to work on spect. DAQ problem
  • 155955 09/25/05 22:49 pking Left arm dipole tripped.
  • 155954 09/25/05 22:47 pking Moved target to empty position.
  • 155951 09/25/05 22:06 pking Moller quads can't be turned back on.
  • 155950 09/25/05 21:46 rsholmes Moller quad
  • 155949 09/25/05 21:36 pking Compton run 12302
  • 155946 09/25/05 21:11 pking Target screens
  • 155945 09/25/05 21:10 Moffit Correction... flipping setpoints did not flip position difference signs
  • 155944 09/25/05 21:09 rsholmes Hall A tools
  • 155943 09/25/05 21:01 Moffit Helicity magnet slopes
  • 155942 09/25/05 20:34 Moffit Summary of helicity magnet scans
  • 155939 09/25/05 20:23 Moffit 5374: Baseline with helicity magnets off
  • 155936 09/25/05 20:06 Moffit 5372 not production. Feedback aborted
  • 155931 09/25/05 19:34 pking First few minutes of Compton run 12300 had IHWP out, the rest was IHWP in.
  • 155930 09/25/05 19:25 pking Compton run 12300.
  • 155925 09/25/05 18:31 pking Compton run 12299
  • 155918 09/25/05 17:40 Moffit HRS DAQ: ROCs 2-4 refuse to load bootscripts
  • 155917 09/25/05 17:25 pking Results of Compton run 12298
  • 155914 09/25/05 16:22 Armstrong Day Shift Summary
  • 155912 09/25/05 16:17 Armstrong BTA will not submit
  • 155911 09/25/05 16:02 Armstrong scanner plots
  • 155907 09/25/05 15:37 Armstrong Compton result run 12297
  • 155905 09/25/05 15:10 Armstrong more target screens
  • 155904 09/25/05 15:06 Armstrong Target screens
  • 155903 09/25/05 14:47 G.M. Urciuol Hall A Tools
  • 155898 09/25/05 12:45 Armstrong Compton cavity not locked
  • 155896 09/25/05 12:34 Armstrong IHWP out
  • 155894 09/25/05 11:41 Armstrong RHWP scan with Hall C off
  • 155888 09/25/05 10:37 Armstrong start RHWP source scan study
  • 155884 09/25/05 09:21 Armstrong correct run comment for 5354
  • 155879 09/25/05 08:02 A. Saha Shift Summary
  • 155875 09/25/05 06:15 Saha IHWP IN
  • 155872 09/25/05 05:12 A. Saha Compton crash after 12293
  • 155871 09/25/05 05:11 A. Saha Compton 12293
  • 155868 09/25/05 04:29 Moffit Helicity corrector magnet slopes (5221-5241)
  • 155863 09/25/05 03:12 A. Saha Compton Run 12293
  • 155862 09/25/05 02:22 Moffit Profile scans at different HVs (5346 and 5347)
  • 155853 09/25/05 00:24 A. Saha Compton Run 12292
  • 155850 09/25/05 00:01 X. Zheng swing shift summary
  • Saturday

  • 155848 09/24/05 23:47 Moffit Left arm dither-corrected detector asymmetry okay.
  • 155846 09/24/05 23:18 X. Zheng beam for profile scan 5343
  • 155842 09/24/05 22:46 A. Deur beam modulation plots look weird ?
  • 155841 09/24/05 22:11 X. Zheng Compton 12291 Pb=80.6%
  • 155840 09/24/05 22:04 X. Zheng Beam IHWP retracted
  • 155835 09/24/05 20:35 lkaufman WAC Summary for Slug 112
  • 155834 09/24/05 20:31 lkaufman WAC Summary for Slug 111
  • 155833 09/24/05 20:13 X. Zheng Compton background rates higher
  • 155830 09/24/05 20:00 A. Deur target charts
  • 155829 09/24/05 19:59 X. Zheng Hall A Tools screen captured
  • 155824 09/24/05 17:33 A. Deur CHL flow fluctuation
  • 155823 09/24/05 17:27 X. Zheng unstable beam
  • 155820 09/24/05 16:46 X. Zheng Compton 12289
  • 155817 09/24/05 16:18 X. Zheng Beam IHWP inserted
  • 155814 09/24/05 16:01 M. Finn Day shift summary
  • 155812 09/24/05 15:50 Yi Qiang Hall A TOol Screen
  • 155811 09/24/05 15:32 Xiaohui Zhan Polarity 88.1% Compton run 12288
  • 155810 09/24/05 15:24 Xiaohui Zhan Polarity 85.8% Compton run 12287
  • 155803 09/24/05 12:32 Xiaohui Zhan Target tools
  • 155800 09/24/05 11:33 M. Finn IHWP is now OUT
  • 155797 09/24/05 10:42 M. Finn compton roc crash
  • 155789 09/24/05 08:00 kfuoti Owl shift summary
  • 155784 09/24/05 07:05 Adaq Hall A Tools
  • 155783 09/24/05 07:03 kfuoti Target screens
  • 155779 09/24/05 06:45 kfuoti IHWP IN
  • 155762 09/24/05 01:59 kfuoti IHWP out
  • 155755 09/24/05 01:10 kfuoti Profile Scan - Run 5319
  • 155750 09/24/05 00:10 X. Zheng Swing shift summary
  • 155749 09/24/05 00:08 A. Deur VESDA alarm on the other HRS
  • 155747 09/24/05 00:02 lkaufman Large Aq scans for detector linearity (Runs 5315 and 5316)
  • Friday

  • 155746 09/23/05 23:41 R. Snyder feedback & dithering turned back ON for production running
  • 155744 09/23/05 23:38 lkaufman Large Aq scans for detector linearity (Runs 5315 and 5316)
  • 155740 09/23/05 23:00 X. Zheng VESDA alarm
  • 155739 09/23/05 22:56 Jack Right Arm Shield House VESDA
  • 155736 09/23/05 22:14 R. Snyder Runlist for Source Scans with Hall C off
  • 155733 09/23/05 21:32 X. Zheng more plots for cavity power
  • 155732 09/23/05 21:30 X. Zheng Compton epics restart
  • 155731 09/23/05 21:27 X. Zheng Compton cavity power low?
  • 155726 09/23/05 20:57 X. Zheng spot++
  • 155720 09/23/05 19:46 H. Benaoum Regression & Dithering Slopes for Detectors vs. Slug (1-110) without slugs 1,2,3,39,40 & 41.
  • 155719 09/23/05 19:24 Moffit Tiefenback Energy vs slug (1-110)
  • 155718 09/23/05 16:46 H. Benaoum Differences for Batteries versus Slug Number (slug 1 till 110 ).
  • 155717 09/23/05 16:18 A. Deur Target in 20cm He position
  • 155715 09/23/05 16:01 H. Benaoum Day Shift Summary
  • 155714 09/23/05 13:47 lkaufman Dithering and feedback off
  • 155710 09/23/05 12:06 lkaufman WAC Daily Summary for Sept. 22
  • 155707 09/23/05 11:47 lkaufman WAC Summary for Slug 110
  • 155706 09/23/05 11:44 lkaufman WAC Summary for Slug 109
  • 155705 09/23/05 11:35 H. Benaoum Beam Polarization Measured by Compton ( September 12 till 23 ).
  • 155698 09/23/05 09:01 kfuoti Scanner access
  • 155695 09/23/05 08:15 fansler Techs/late entry
  • 155693 09/23/05 08:06 G.M. Urciuol Owl shift report
  • 155691 09/23/05 06:17 G.M. Urciuol Hall A Tools screen
  • 155690 09/23/05 06:14 kfuoti target screens
  • 155680 09/23/05 04:47 kfuoti Right septum heating
  • 155679 09/23/05 04:40 kfuoti Left septum quench - again
  • 155674 09/23/05 02:19 kfuoti Left Septum Quench
  • 155664 09/23/05 00:44 G.M. Urciuol IHWP
  • 155658 09/23/05 00:18 Moffit Re: 5293 profile scan - left arm
  • 155657 09/23/05 00:03 Ulmer Swing Shift Summary
  • Thursday

  • 155654 09/22/05 23:42 Ulmer Compton misses a beat
  • 155652 09/22/05 23:29 Ulmer Compton cavity troubles?
  • 155649 09/22/05 22:35 Ulmer Compton Run 12273
  • 155648 09/22/05 22:07 Ulmer Profile Scan Results
  • 155643 09/22/05 20:41 Ulmer Target Charts
  • 155642 09/22/05 20:38 Ulmer Hall A Tools
  • 155639 09/22/05 19:58 Ulmer IHWP=OUT
  • 155638 09/22/05 19:51 Ulmer Beam position adjusted
  • 155637 09/22/05 19:48 Ulmer Compton Run 12272
  • 155634 09/22/05 19:09 Ulmer Compton Run 12271
  • 155633 09/22/05 18:58 R. Snyder 4A & 4B position differences since Hall A laser lock fix
  • 155632 09/22/05 18:54 R. Snyder 4A & 4B position differences since labor day
  • 155627 09/22/05 18:12 E.Chudakov Spot
  • 155623 09/22/05 17:10 R. Snyder 4A & 4B position differences through 108 slugs
  • 155622 09/22/05 17:06 Armstrong the tour today
  • 155621 09/22/05 16:48 Moffit Integrated aluminum fraction (from run 5207)
  • 155618 09/22/05 16:04 X. Zheng Day shift summary
  • 155616 09/22/05 15:48 Armstrong move target; start to unpark
  • 155613 09/22/05 15:31 spiegel Q1 Danfysik power supply
  • 155612 09/22/05 15:12 H. Benaoum Cavity Monitors Position Differences versus Slug Number ( till slug 108).
  • 155609 09/22/05 14:27 Moffit HAPPEX DAQ max event limit changed back to 100k
  • 155606 09/22/05 13:40 Moffit left arm detector pedestal width smaller.
  • 155603 09/22/05 13:14 Armstrong coulombometer update
  • 155602 09/22/05 13:05 Armstrong Left Septum trip redux
  • 155601 09/22/05 12:42 H. Benaoum Raw, Regressed, Dithering Corrected and Regression Corrected Detectors Normalized Asymmetries versus Slug Number (1-108) without slug 39, 40 and 41.
  • 155600 09/22/05 12:23 Armstrong septum reset OK
  • 155599 09/22/05 11:45 Armstrong left septum trip.
  • 155598 09/22/05 10:40 Moffit HAPPEX DAQ: max event limit to 800k
  • 155595 09/22/05 10:20 H. Benaoum Raw and Regressed Detectors Normalized Asymmetries versus Slug Number (77-108 ) after transverse polarization flip.
  • 155594 09/22/05 09:46 X. Zheng Hall status changed to restricted access
  • 155591 09/22/05 09:29 Armstrong Park target
  • 155590 09/22/05 08:54 E.Chudakov Restricted Access now
  • 155586 09/22/05 08:02 G.M. Urciuol Owl shift report
  • 155583 09/22/05 07:14 kfuoti target screens
  • 155582 09/22/05 07:13 kfuoti Tools Screen
  • 155576 09/22/05 04:12 lkaufman ROC 26 reboot
  • 155573 09/22/05 04:00 kfuoti RQ1 trip
  • 155572 09/22/05 03:40 lkaufman WAC Daily Summary for Sept. 21
  • 155563 09/22/05 01:57 lkaufman WAC Summary Slug 108
  • 155562 09/22/05 01:53 lkaufman WAC Summary Slug 107
  • 155561 09/22/05 01:46 lkaufman WAC Summary Slug 106
  • 155560 09/22/05 01:43 lkaufman WAC Summary Slug 105
  • 155549 09/22/05 00:01 Yi Qiang Swing Shift Summary
  • Wednesday

  • 155547 09/21/05 23:35 Yi Qiang Profile Scan Plots
  • 155546 09/21/05 23:28 Yi Qiang Change IHWP to IN
  • 155541 09/21/05 22:55 BERTIN P.Y. target screen captur
  • 155540 09/21/05 22:34 Yi Qiang Hall A Tool Screen
  • 155539 09/21/05 22:21 Yi Qiang Spot Size
  • 155535 09/21/05 22:06 E. Chudakov Left Q1 trip and recovery
  • 155534 09/21/05 20:40 R. Snyder 4A & 4B position differences since M. Poelker's Hall A laser lock fix
  • 155533 09/21/05 20:38 R. Snyder 4A & 4B position differences from labor day up to M. Poelker's Hall A laser lock fix
  • 155532 09/21/05 20:31 R. Snyder 4A & 4B position differences since labor day
  • 155530 09/21/05 20:00 Yi Qiang Left Q1 tripped
  • 155521 09/21/05 17:35 R. Snyder 4A & 4B position differences through 104 slugs
  • 155518 09/21/05 16:06 jpchen day shift summary
  • 155514 09/21/05 15:45 jpchen Compton is back running
  • 155513 09/21/05 15:43 jpchen IHWP OUT
  • 155508 09/21/05 15:25 S. Fuchs target screen
  • 155507 09/21/05 15:25 jpchen magnet screen
  • 155506 09/21/05 15:19 jpchen Compton down since noon
  • 155505 09/21/05 14:53 Moffit Curious PAN fEvLen errors..
  • 155502 09/21/05 14:42 Moffit Access activity.
  • 155497 09/21/05 13:21 jpchen Left Q1 and Right Q1 lost, right Q1 reset ok.
  • 155493 09/21/05 10:28 lkaufman Cavities checked with spectrometer DAQ
  • 155491 09/21/05 10:23 jpchen beam current fluctuation
  • 155488 09/21/05 10:12 jpchen Lost left magnets control. Reboot ioc and reset.
  • 155487 09/21/05 10:05 lkaufman Cavity monitors stopped working - working again!
  • 155483 09/21/05 09:36 suleiman Coulombometer Update
  • 155479 09/21/05 09:20 jpchen Beam current fluctuation
  • 155478 09/21/05 09:15 lkaufman WAC Daily Summary for Sept. 20
  • 155476 09/21/05 09:11 lkaufman Rebooted ROCs and CODA for spectrometer DAQ
  • 155473 09/21/05 08:44 G.M. Urciuol ROCS rebooted
  • 155469 09/21/05 08:18 lkaufman WAC Summary Slug 104
  • 155468 09/21/05 08:15 lkaufman WAC Summary Slug 103
  • 155467 09/21/05 08:14 vasulk Owl Shift Summary
  • 155466 09/21/05 08:13 lkaufman WAC Summary Slug 102
  • 155465 09/21/05 08:09 lkaufman WAC Summary Slug 101
  • 155463 09/21/05 07:37 vasulk Polarization from Compton: 82.3%
  • 155462 09/21/05 07:30 kfuoti IHWP in
  • 155459 09/21/05 06:41 vasulk Compton Run 12258, P = 82.1%
  • 155456 09/21/05 05:50 kfuoti Target Screens
  • 155455 09/21/05 05:48 kfuoti Hall A Tools Screen
  • 155452 09/21/05 04:50 vasulk Polarization from Compton: 82.8%
  • 155449 09/21/05 04:20 vasulk Raster Check
  • 155441 09/21/05 03:11 vasulk Compton: Telnet Problem for ROCS
  • 155440 09/21/05 02:50 kfuoti IHWP out
  • 155433 09/21/05 00:38 vasulk Compton Run 12254, P = -81.1%
  • 155431 09/21/05 00:02 A. Puckett swing shift summary--complete
  • Tuesday

  • 155429 09/20/05 23:45 A. Puckett ioc reboot left HRS magnets
  • 155425 09/20/05 22:38 A. Puckett Nice beam...
  • 155421 09/20/05 22:06 PY BERTIN half wave in
  • 155420 09/20/05 22:02 Suleiman Helicity Magnets Calibration
  • 155373 09/20/05 19:51 A. Puckett tgt screen capture for shift chklist
  • 155371 09/20/05 19:49 A. Puckett tgt screen capture for shift chklist
  • 155370 09/20/05 19:46 A. Puckett Hall A tools screen capture for shift chklist
  • 155366 09/20/05 19:39 A. Puckett helicity magnet test, RHWP scan
  • 155358 09/20/05 16:24 A. Puckett swing shift summary
  • 155356 09/20/05 16:00 X. Zheng day shift summary
  • 155348 09/20/05 14:58 G.M. Urciuol HallA Tools
  • 155347 09/20/05 14:56 X. Zheng Target screen capture
  • 155340 09/20/05 13:51 H. Benoum Beam Polarization Measured by Compton Polarimeter from 12th till today
  • 155339 09/20/05 13:42 H. Benaoum Regression & Dithering Slopes for Detectors vs. Slug (1-100) without slugs 1,2,3,39,40 & 41.
  • 155335 09/20/05 13:16 X. Zheng Beam IHWP changed to OUT
  • 155334 09/20/05 13:15 H. Benaoum Raw, Regressed, Dithering Corrected and Regression Corrected Detectors Normalized Asymmetries versus Slug Number (1-100) without slug 39, 40 and 41.
  • 155332 09/20/05 13:07 H. Benaoum Raw and Regressed Detectors Normalized Asymmetries versus Slug Number (77-100 ) after transverse polarization flip.
  • 155331 09/20/05 13:05 H. Benaoum Raw and Regressed Detectors Normalized Asymmetries versus Slug Number (77-100 ) after transverse polarization flip.
  • 155320 09/20/05 11:00 X. Zheng MCC IOC reboot
  • 155317 09/20/05 10:14 X. Zheng Work on Hall C steering
  • 155314 09/20/05 10:10 X. Zheng Target moved to 20cm LHe
  • 155313 09/20/05 09:26 X. Zheng Target moved to 20cm Al dummy
  • 155308 09/20/05 08:05 vasulk Owl Shift Summary
  • 155304 09/20/05 07:20 Moffit Bleedthrough measurement during run 5204
  • 155303 09/20/05 06:52 vasulk Compton Run 12247 Junk!
  • 155302 09/20/05 06:50 Moffit Run 5203: Profile scan
  • 155301 09/20/05 06:40 Moffit IHWP IN!
  • 155296 09/20/05 05:46 lkaufman WAC Daily Summary for Sept. 19
  • 155295 09/20/05 05:31 lkaufman WAC Summary Slug 100
  • 155292 09/20/05 05:28 lkaufman WAC Summary Slug 99
  • 155291 09/20/05 05:25 lkaufman WAC Summary Slug 98
  • 155290 09/20/05 05:23 lkaufman WAC Summary Slug 97
  • 155286 09/20/05 04:43 Moffit spot++ captured!
  • 155283 09/20/05 04:27 vasulk Compton S/B: 8.5
  • 155282 09/20/05 04:15 R. Snyder 4A & 4B position differences since Ryan's source change up to Hall laser C lock failure
  • 155281 09/20/05 04:13 Moffit Target screen captured!
  • 155280 09/20/05 04:11 vasulk Hall A Tools Screen
  • 155279 09/20/05 04:00 vasulk Lost Left Arm Magent IOC
  • 155278 09/20/05 03:44 Moffit Run 5199: ROC26 rebooted itself
  • 155276 09/20/05 03:22 R. Snyder 4A & 4B position differences since labor day up to Ryan's source change
  • 155273 09/20/05 02:15 Moffit IHWP OUT!
  • 155270 09/20/05 02:04 vasulk Compton Run 12243
  • 155265 09/20/05 00:30 Moffit spot++ captured!
  • 155261 09/20/05 00:02 Ulmer Swing shift summary
  • 155260 09/20/05 00:01 Ulmer Compton Run 12242
  • Monday

  • 155259 09/19/05 23:33 Ulmer Compton Run 12241 with bad y-slope
  • 155256 09/19/05 22:33 Yi Qiang Profile Scan 5913 Plots
  • 155251 09/19/05 21:37 Ulmer Compton Run 12237 - try again
  • 155250 09/19/05 21:31 Ulmer Compton Run 12240
  • 155249 09/19/05 21:29 Ulmer Compton Run 12237
  • 155248 09/19/05 21:25 Ulmer Run 5191 - beam steering
  • 155247 09/19/05 21:19 Yi Qiang Change IHWP to IN
  • 155245 09/19/05 21:18 Ulmer Compton Run 12237 vertical position
  • 155243 09/19/05 21:09 pking Spot++ after the Compton chicane correction
  • 155240 09/19/05 20:28 R. Snyder WAC Summary Slug 94
  • 155237 09/19/05 20:07 R. Snyder WAC Summary Slug 93
  • 155234 09/19/05 18:50 R. Snyder WAC Sumary Slug 92
  • 155233 09/19/05 18:49 Ulmer Compton Run 12235 IHWP
  • 155232 09/19/05 18:47 P. King Target charts
  • 155231 09/19/05 18:45 R. Snyder WAC Summary Slug 91
  • 155228 09/19/05 18:10 Ulmer Hall A Tools
  • 155227 09/19/05 17:58 Ulmer HV trip
  • 155223 09/19/05 16:31 Yi Qiang Change IHWP to OUT
  • 155221 09/19/05 16:06 Yi Qiang Profile Scan plots of 5185
  • 155220 09/19/05 16:01 X. Zheng day shift summary
  • 155218 09/19/05 15:53 H. Benaoum -83.9% Beam Polarization Measured by Compton (run 12233)
  • 155207 09/19/05 14:16 X. Zheng Hall A Tools Screen & Septa strip charts
  • 155206 09/19/05 14:15 S. Fuchs target screen
  • 155205 09/19/05 14:05 Suleiman Feedback for Hall-C is OFF - Back to Hall B laser
  • 155204 09/19/05 13:55 H. Benaoum Raw and Regressed Detectors Normalized Asymmetries versus Slug Number (77-92) after polarization flip
  • 155203 09/19/05 13:51 H. Benaoum Raw, Regressed, Dithering Corrected and Regression Corrected Detectors Normalized Asymmetries versus Slug Number (1-92) without slugs 39,40 and 41
  • 155200 09/19/05 12:52 Suleiman feedback for hall-C is back ON
  • 155197 09/19/05 12:06 R. Feuerbach Q2 measurement
  • 155180 09/19/05 08:55 spiegel Q1 failure 100A max. ramp-up
  • 155177 09/19/05 08:03 Moffit Owl Shift Summary
  • 155175 09/19/05 07:59 vasulk Inserted Beam Half-wave Plate
  • 155168 09/19/05 05:42 vasulk Target Screens Shot
  • 155167 09/19/05 05:40 vasulk Hall A Tools Screen Shot
  • 155166 09/19/05 05:30 vasulk Periodic drops in Right Q1 He Level
  • 155164 09/19/05 05:14 vasulk Raster Check
  • 155158 09/19/05 03:19 vasulk IHWP Retracted
  • 155152 09/19/05 01:52 vasulk Spot/Raster Check
  • 155148 09/19/05 00:31 Moffit Hall-C using Hall-B laser. Feedback on hall-C off
  • 155143 09/19/05 00:01 M. Finn Swing shift summary
  • Sunday

  • 155138 09/18/05 22:22 M. Finn IHWP is now IN
  • 155136 09/18/05 21:36 kfuoti Target Screens
  • 155135 09/18/05 21:34 kfuoti Hall A Tools Screen
  • 155133 09/18/05 20:58 Adaq IOC crash on left HRS
  • 155123 09/18/05 16:44 M. Finn IHWP changed to out position
  • 155122 09/18/05 16:43 M. Finn run 5160 aborted
  • 155117 09/18/05 16:02 H. Benaoum Day Shift Summary
  • 155115 09/18/05 16:01 H. Benaoum Beam Polarization measured by Compton Polarimeter from 12th-18th September
  • 155114 09/18/05 15:55 R. Snyder Hall A Target Screens
  • 155113 09/18/05 15:46 R. Snyder Profile Scan --Run 5158
  • 155108 09/18/05 14:53 R. Snyder WAC Summary for Slug 90
  • 155107 09/18/05 14:45 H. Benaoum adaql1 keyboard not working
  • 155106 09/18/05 14:33 R. Snyder WAC Summary Slug 89
  • 155103 09/18/05 14:11 H. Benaoum No Coincidence rates during the end of compton run 12215
  • 155102 09/18/05 14:05 R. Snyder WAC Summary Slug 88
  • 155097 09/18/05 12:35 H. Benaoum Compton Quarter Wave Plate Scan
  • 155096 09/18/05 12:10 R. Snyder WAC Summary Slug 87
  • 155093 09/18/05 11:39 R. Snyder new PC Voltages for IHWP.IN
  • 155088 09/18/05 10:42 R. Snyder IHWP IN before run 5151
  • 155083 09/18/05 09:18 H. Benaoum 83.2% Beam Polarization measured by compton (run 12212)
  • 155080 09/18/05 08:21 H. Benaoum Compton Rocs problem fixed
  • 155079 09/18/05 08:06 Moffit Owl Shift Summary
  • 155075 09/18/05 07:46 Moffit Compton: ROCs frozen.
  • 155072 09/18/05 06:21 Moffit Run 5145: profile scan (production momentum).
  • 155069 09/18/05 05:57 Apar target screen shot
  • 155067 09/18/05 05:53 Moffit IHWP OUT!
  • 155064 09/18/05 05:38 Moffit Back to production p0. Right Q1 set by hand
  • 155063 09/18/05 05:29 Moffit -10MeV profile scan
  • 155061 09/18/05 05:17 Moffit Re: +0.5% right arm splash back
  • 155059 09/18/05 05:11 Moffit +0.5% profile scan
  • 155057 09/18/05 05:01 Moffit Right Q1 trip
  • 155055 09/18/05 04:51 Moffit +1% profile scan: Still too far away.
  • 155052 09/18/05 04:32 Moffit Techs: Right Q1 wont lock to p0 set
  • 155051 09/18/05 04:30 Moffit +2% profile scan. No good.
  • 155048 09/18/05 03:57 Moffit 0%: Profile scan.
  • 155042 09/18/05 02:39 Moffit Compton: End of run 12209
  • 155040 09/18/05 01:54 Moffit spot++ captured!
  • 155034 09/18/05 00:05 Yi Qiang Swing shift summary
  • Saturday

  • 155031 09/17/05 23:59 Yi Qiang change IHWP to IN
  • 155027 09/17/05 23:18 Yi Qiang ROC1 lost connection
  • 155024 09/17/05 21:58 Yi Qiang Screen Capture
  • 155023 09/17/05 21:56 Yi Qiang compton polarity 85.6% from run 12205 ended by coda crash
  • 155018 09/17/05 20:34 Xiaohui Zhan Hall A tool screen
  • 155017 09/17/05 20:32 Xiaohui Zhan Hall A tool screen
  • 155016 09/17/05 20:29 Xiaohui Zhan Hall A tool screen
  • 155015 09/17/05 20:25 Yi Qiang Polarity 82.2% from comptom run 12203
  • 155014 09/17/05 19:55 Xiaohui Zhan IHWP out
  • 155009 09/17/05 17:58 Yi Qiang Recover Left HRS control
  • 155004 09/17/05 16:39 Yi Qiang Profile Scan # 5126
  • 154999 09/17/05 16:01 Mike Finn Day shift summary
  • 154995 09/17/05 15:06 H. Benaoum Target Screen
  • 154994 09/17/05 15:06 M. Finn Tool screen capture
  • 154992 09/17/05 14:15 H. Benaoum Spot++
  • 154982 09/17/05 11:25 H. Benaoum Raw & Regressed Normalized Detector Asymmetries versus Slug (77-86)
  • 154981 09/17/05 11:11 H. Benaoum Raw, Regressed, Dithering Corrected and Regression Corrected Detectors Normalized Asymmetries versus Slug Number (1-86) without slug 39,40 and 41.
  • 154980 09/17/05 11:11 M. Finn snapshot of Hall A Tools
  • 154979 09/17/05 11:08 M. Finn septum magnet chart profile
  • 154978 09/17/05 11:07 M. Finn reset septum magnets
  • 154976 09/17/05 10:42 Apar changed IHWP to IN position
  • 154972 09/17/05 09:27 H. Benaoum Compton Polarization 82.2% from run 12196
  • 154969 09/17/05 08:02 Moffit Owl Shift Summary
  • 154961 09/17/05 05:49 Moffit IHWP OUT!
  • 154958 09/17/05 05:44 R. Snyder New IHWP OUT Source Setting
  • 154957 09/17/05 05:23 R. Snyder New PC Voltages for IHWP OUT-- flipper file updated
  • 154956 09/17/05 05:20 R. Snyder Hall A Target Screens
  • 154955 09/17/05 05:18 R. Snyder new PITA slope for IHWP OUT
  • 154954 09/17/05 05:18 Moffit Hall-A Tools screen captured!
  • 154953 09/17/05 05:17 R. Snyder Runlist from Source adjustment
  • 154951 09/17/05 04:59 Moffit spot++ captured!
  • 154949 09/17/05 04:54 R. Snyder WAC Summary Slug 86
  • 154946 09/17/05 04:47 R. Snyder WAC Summary Slug 85
  • 154941 09/17/05 02:42 Moffit 5107, 5108, 5109 are parityscans.
  • 154940 09/17/05 02:02 Moffit IHWP IN!
  • Friday

  • 154928 09/16/05 23:58 Aniol swing shift summary
  • 154927 09/16/05 23:38 Aniol compton roc5 crashes/restarted
  • 154924 09/16/05 23:05 Aniol compton run 12192 started
  • 154923 09/16/05 22:59 Moffit Left arm Profile scans, with and without right septum
  • 154920 09/16/05 22:16 Aniol compton photon HV trips off
  • 154919 09/16/05 21:50 Aniol start compton run 12191
  • 154918 09/16/05 21:41 Yi Qiang Set IHWP to OUT
  • 154916 09/16/05 21:36 Aniol end compton run 12190
  • 154914 09/16/05 20:38 Aniol septa temps
  • 154912 09/16/05 20:23 Aniol septa strip charts for quench
  • 154911 09/16/05 20:09 Aniol beam off to cool down left septum
  • 154908 09/16/05 20:03 Aniol left septum quenched
  • 154904 09/16/05 18:58 Xiaohui Zhan Target Screen
  • 154903 09/16/05 18:44 Aniol start compton 12190
  • 154902 09/16/05 18:42 Aniol back in production with 2 spec
  • 154900 09/16/05 18:35 Yi Qiang Hall A Tool Screen
  • 154899 09/16/05 18:30 Yi Qiang Spot++
  • 154896 09/16/05 18:25 Aniol beam starts into hall
  • 154895 09/16/05 18:10 Aniol go to power permit
  • 154893 09/16/05 17:52 Aniol shedule for tech on call
  • 154892 09/16/05 17:51 R. Snyder 4A & 4B position differences through 84 slugs
  • 154891 09/16/05 17:48 Aniol turning on all magnets
  • 154890 09/16/05 17:43 R. Snyder WAC Summary Slug 84
  • 154889 09/16/05 17:35 fansler techs
  • 154888 09/16/05 17:23 R. Snyder WAC Summary Slug 83
  • 154885 09/16/05 16:17 Aniol move target to empty
  • 154882 09/16/05 16:10 H. Benaoum Day Shift Summary
  • 154876 09/16/05 14:34 fansler Right Q1 status
  • 154873 09/16/05 13:59 Fuchs target screen
  • 154864 09/16/05 10:01 H. Benaoum Compton Quarter Wave Plate Scan
  • 154859 09/16/05 08:08 lkaufman Owl Shift Summary
  • 154855 09/16/05 06:48 kfuoti Target screens
  • 154854 09/16/05 06:46 kfuoti Hall A Tools Screen
  • 154849 09/16/05 04:57 lkaufman No beam until day shift
  • 154848 09/16/05 04:32 stevensm tech/right Q1 won't ramp above 100A
  • 154846 09/16/05 04:22 lkaufman Hall A Laser not locking
  • 154844 09/16/05 04:06 lkaufman R Q1 still needs work, so just LHRS for now
  • 154839 09/16/05 02:17 lkaufman Still in controlled access
  • 154836 09/16/05 00:55 lkaufman Controlled Access for RQ1
  • 154833 09/16/05 00:02 R. Snyder Swing Shift Summary
  • Thursday

  • 154832 09/15/05 23:11 R. Snyder Q1 trouble
  • 154831 09/15/05 22:18 R. Snyder Rt. Q1 tripped; need controlled acess
  • 154830 09/15/05 21:44 R. Snyder WAC Summary Slug 82
  • 154828 09/15/05 20:31 R. Snyder Rt. Septum tripped at 7.70 K
  • 154825 09/15/05 20:06 Aniol target chart
  • 154824 09/15/05 20:03 Ulmer Hall A Tools
  • 154823 09/15/05 19:44 R. Snyder Profile Scan Run 5077 (Left Arm Only)
  • 154821 09/15/05 19:17 R. Snyder WAC Summary Slug 81
  • 154817 09/15/05 17:59 R. Snyder Rt. Septum tripped at 7.73 K
  • 154814 09/15/05 16:45 H. Benaoum Compton Polarization from run 12186
  • 154813 09/15/05 16:43 R. Snyder IHWP flipped IN before run 5075
  • 154808 09/15/05 16:29 H. Benaoum Raw and Regressed Detectors Normalized Asymmetries versus Slug Number (77-80).
  • 154807 09/15/05 16:25 H. Benaoum Raw, Regressed, Dithering Corrected and Regression Corrected Detectors Normalized Asymmetries versus Slug Number (1-80) without slug 39,40 and 41.
  • 154805 09/15/05 16:00 Saha Shift Summary
  • 154804 09/15/05 15:54 S. Bailey Helicity Magnet Test: Hall C Results
  • 154801 09/15/05 15:18 R. Snyder WAC Summary Slug 80
  • 154800 09/15/05 15:11 Reitz EP Beam Energy Measurement
  • 154797 09/15/05 15:01 Bosted Raster
  • 154794 09/15/05 14:37 Bosted Target moved back in beam, unparked
  • 154793 09/15/05 12:13 Moffit Quick online result of helicity correct magnet runs
  • 154792 09/15/05 12:10 nanda compton data
  • 154791 09/15/05 11:15 Armstrong Coulombometer update
  • 154790 09/15/05 10:36 Bosted Target parked
  • 154789 09/15/05 10:14 Bosted Target to MT
  • 154788 09/15/05 10:10 Saha Septum Quench Studies
  • 154781 09/15/05 08:18 Suleiman Helicity Magnets are cleared to run
  • 154780 09/15/05 08:03 lkaufman Owl Shift Summary
  • 154776 09/15/05 06:59 kfuoti Profile Scan: Run 5067
  • 154775 09/15/05 06:54 lkaufman Compton Run 12180 Polarization measurement
  • 154774 09/15/05 06:46 lkaufman IHWP flippered to OUT
  • 154767 09/15/05 04:57 kfuoti target screens
  • 154766 09/15/05 04:55 kfuoti Hall A tools screen
  • 154759 09/15/05 02:04 kfuoti IHWP in before run 5062
  • 154756 09/15/05 01:40 lkaufman Spect. DAQ restarted
  • 154752 09/15/05 00:15 R. Snyder Swing Shift Summary
  • 154751 09/15/05 00:08 lkaufman re-prompted runs 4989-5049
  • Wednesday

  • 154746 09/14/05 22:13 R. Snyder IHWP flipped to OUT before run 5057
  • 154745 09/14/05 22:06 R. Snyder Dithering was turned off for last half of Slug 83
  • 154742 09/14/05 21:07 Aniol compton run 12177
  • 154733 09/14/05 18:20 R. Snyder spot++ raster check
  • 154730 09/14/05 16:54 H. Benaoum Compton Quarter Wave Plate Scan
  • 154727 09/14/05 16:02 R. Michaels day summary
  • 154726 09/14/05 16:02 Aniol target charts
  • 154725 09/14/05 16:00 Adaq tools
  • 154723 09/14/05 13:53 S. Bailey Injector Helicity Magnet Scans Summary (corrected G0 run numbers & readable table)
  • 154722 09/14/05 13:46 S. Bailey Injector Helicity Magnet Scans Summary (corrected G0 run numbers)
  • 154721 09/14/05 13:11 R. Michaels Beam down
  • 154717 09/14/05 11:44 lkaufman Changed lobeam in .db back to 25000
  • 154714 09/14/05 11:35 Suleiman Run Plan
  • 154713 09/14/05 11:05 S. Bailey Injector Helicity Magnet Scans Summary (again)
  • 154712 09/14/05 10:58 S. Bailey Injector Helicity Magnet Scans Summary
  • 154699 09/14/05 09:29 R. Michaels beam very stable
  • 154693 09/14/05 08:24 R. Michaels cavity check -- good
  • 154691 09/14/05 08:01 kfuoti OWL shift summary
  • 154687 09/14/05 07:10 kfuoti Scan Comparison
  • 154682 09/14/05 05:51 vasulk Profile Scan, Run # 5038
  • 154681 09/14/05 05:45 vasulk Beam Half-wave Plate Retracted
  • 154676 09/14/05 04:43 kfuoti target screens
  • 154675 09/14/05 04:39 kfuoti Hall A tools screen capture
  • 154670 09/14/05 02:55 vasulk Empty Histograms
  • 154667 09/14/05 02:11 kfuoti spot++
  • 154662 09/14/05 00:43 vasulk Inserted Beam Half-Wave Plate
  • 154659 09/14/05 00:08 R. Snyder 12x, Aq, and Double Differences through 79 slugs
  • 154658 09/14/05 00:05 R. Snyder Swing Shift Summary
  • Tuesday

  • 154654 09/13/05 23:18 Aniol tools screen
  • 154653 09/13/05 22:57 pking Reboot of CODA and the ROCs
  • 154645 09/13/05 21:06 R. Snyder 4A & 4B position differences thorugh 79 slugs
  • 154644 09/13/05 20:54 R. Snyder WAC Summary Slug 79
  • 154643 09/13/05 20:31 R. Snyder WAC Summary Slug 78
  • 154637 09/13/05 19:29 R. Snyder WAC Summary Slug 77
  • 154636 09/13/05 19:29 pking Target screen capture
  • 154631 09/13/05 16:23 Suleiman Heleicty Magnet Test is moved to Wednesday
  • 154629 09/13/05 16:08 R. Snyder IHWP flipped OUT before run 5026
  • 154628 09/13/05 16:03 H. Benaoum Day Shift Summary
  • 154625 09/13/05 15:17 Suleiman Transverse Polarization Flip
  • 154624 09/13/05 15:16 Suleiman Helicity Magnet Test Update
  • 154622 09/13/05 14:33 Bosted Target unparked
  • 154621 09/13/05 14:29 R. Snyder spot++ raster check
  • 154618 09/13/05 14:16 Bosted Target moved to He
  • 154615 09/13/05 12:54 H. Benaoum Position Difference along 4ax smaller compared to before but still big
  • 154611 09/13/05 12:32 Moffit Access... det1 pedestal changed..
  • 154606 09/13/05 10:25 Moffit left det pedestal width increased during 5014
  • 154605 09/13/05 09:54 H. Benaoum Compton Quarter Wave Plate Scan
  • 154604 09/13/05 09:40 Bosted He target parked
  • 154602 09/13/05 09:27 Bosted Target to Empty, began parking
  • 154601 09/13/05 09:26 Moffit Left detector pedestal quite large..
  • 154599 09/13/05 09:15 Moffit Fix to feedback code..
  • 154593 09/13/05 08:15 lkaufman Owl Shift Summary
  • 154592 09/13/05 07:51 lkaufman Feedback not updating...again
  • 154587 09/13/05 06:03 lkaufman Run 5014 and 5015 are production runs
  • 154584 09/13/05 05:58 lkaufman spectrometer DAQ beg. and end of run working again
  • 154581 09/13/05 05:37 vasulk Profile Scan, Run# 5013
  • 154580 09/13/05 05:23 lkaufman spot++
  • 154579 09/13/05 05:22 lkaufman Reboot Spectrometer DAQ and ROCs
  • 154577 09/13/05 05:17 vasulk Insert Half-Wave Plate
  • 154570 09/13/05 03:12 vasulk Feedback Not Updating
  • 154567 09/13/05 02:52 lkaufman Target Screen Captured!
  • 154566 09/13/05 02:50 lkaufman Hall A Tools Captured!
  • 154563 09/13/05 02:09 lkaufman Late Entry: Beam back at 01:17
  • 154557 09/13/05 00:48 lkaufman Hall C had to reboot IOC affects all viewers, so no beam
  • 154556 09/13/05 00:40 lkaufman Changed source parameters for IHWP OUT
  • 154550 09/13/05 00:04 P. King Swing shift summary
  • Monday

  • 154549 09/12/05 23:14 pking Begin Compton run 12154.
  • 154547 09/12/05 23:11 Yi Qiang Put Half-Wave Plate OUT
  • 154545 09/12/05 23:01 pking End Compton run 12152; begin Compton run 12153.
  • 154542 09/12/05 22:10 pking End Compton run 12151; begin Compton run 12152.
  • 154539 09/12/05 21:54 pking spot++
  • 154538 09/12/05 21:50 Yi Qiang Hall A tools screen
  • 154535 09/12/05 21:41 pking Target charts.
  • 154534 09/12/05 21:33 pking Hall A laser is working okay so far.
  • 154533 09/12/05 21:32 pking Beam delivery will continue overnight as long as the Hall A laser works.
  • 154532 09/12/05 21:10 pking Back to normal running; Hall A laser is losing its lock.
  • 154529 09/12/05 20:47 H. Benaoum Beam Polarization 83.7% Measured by Compton (run12148)
  • 154528 09/12/05 20:44 pking Beam came back at 19:13; there are lots of beam trips.
  • 154527 09/12/05 20:30 pking Problems ending Compton run 12148; begin Compton run 12151.
  • 154520 09/12/05 19:06 pking Riad has changed the target PID settings
  • 154519 09/12/05 19:02 pking Beam taken to check the Hall C viewers.
  • 154518 09/12/05 19:02 pking Reduced beam current to keep right septum temp down.
  • 154515 09/12/05 18:17 H. Benaoum Compton after retune and vertical Scan
  • 154514 09/12/05 17:59 pking Beam modulation is now enabled.
  • 154513 09/12/05 17:43 pking Increased beam current to 40 uA before run 4998.
  • 154511 09/12/05 17:32 pking Moved helium target into beam.
  • 154509 09/12/05 16:37 pking Moved to empty target.
  • 154508 09/12/05 16:30 nanda compton tune
  • 154506 09/12/05 16:22 suleiman Heater PID
  • 154503 09/12/05 15:15 R. Snyder spot++ raster check
  • 154500 09/12/05 13:33 lkaufman Closed JT valve a bit
  • 154499 09/12/05 13:25 Kumar Target Charts
  • 154494 09/12/05 10:44 nanda compton setup
  • 154493 09/12/05 10:35 nanda compton prolems fixed
  • 154492 09/12/05 09:28 Moffit Incremented slugnumber to 77.
  • Sunday

  • 154483 09/11/05 15:24 Moffit Detector LEDs on during pedestal run 4991
  • 154476 09/11/05 14:09 Moffit Pedestal runs... Detector and Scanner HV on.
  • 154474 09/11/05 12:48 R. Michaels Sieve slits taken out
  • 154473 09/11/05 11:46 R. Michaels controlled access work
  • 154472 09/11/05 03:15 E.Chudakov End of Moller studies
  • 154470 09/11/05 00:03 E.Chudakov Starting Moller measurements
  • Saturday

  • 154468 09/10/05 21:22 Kumar Beam studies done, Moller continues
  • 154465 09/10/05 20:21 lkaufman Helicity plugged in at ISB
  • 154459 09/10/05 19:28 R. Snyder new IA and PITA slopes, new PC Voltages
  • 154458 09/10/05 19:13 lkaufman real-time helicity unplugged in ISB, PC off
  • 154455 09/10/05 18:56 R. Snyder Hall A Tools Screen Capture
  • 154454 09/10/05 18:30 lkaufman Set HV on HAPPEX detectors
  • 154453 09/10/05 18:28 R. Snyder Runlist for Beam Studies on 9/10
  • 154425 09/10/05 14:26 R. Michaels status, run plan
  • 154414 09/10/05 13:03 Kumar raster on
  • 154411 09/10/05 12:59 Kumar Moller done, start beam systematic studies
  • 154403 09/10/05 12:07 E.Chudakov End of Moller measurements
  • 154400 09/10/05 03:15 E.Chudakov Start Moller measurements
  • 154399 09/10/05 02:33 Holmstrom Owl Shift Summary
  • 154398 09/10/05 02:32 Holmstrom Spectrometer checklist
  • 154397 09/10/05 02:31 Holmstrom Sieves still in
  • 154396 09/10/05 02:30 Holmstrom empty target for Moller
  • 154395 09/10/05 02:29 Adaq HV Capture
  • 154381 09/10/05 01:23 H. Benaoum Hall A Tools
  • 154368 09/10/05 00:01 R. Feuerbach Shift summary -- Pointing runs part one
  • Friday

  • 154339 09/09/05 21:01 R. Feuerbach run 2628 was on tantalum
  • 154331 09/09/05 20:28 Yi QiangHa Hall A Tool Screen
  • 154329 09/09/05 20:27 Yi Qiang Target Screen
  • 154324 09/09/05 19:56 lkaufman Lost Left Arm Controls, magnets back
  • 154315 09/09/05 19:03 lkaufman changed lobeam in .db
  • 154299 09/09/05 17:32 nanda compton background rate
  • 154283 09/09/05 15:00 R. Feuerbach sieve slit collimators IN
  • 154282 09/09/05 14:53 Moffit Pedestal runs 4950-4956
  • 154281 09/09/05 14:13 Moffit Septum+HRS setting for +1.5% tantilum elastic
  • 154280 09/09/05 13:09 nanda compton electron detector
  • 154279 09/09/05 12:52 Moffit Hall back to power/beam permit.
  • 154278 09/09/05 12:47 Moffit Controlled Access for Compton
  • 154276 09/09/05 11:26 Moffit Hall A Tools screen captured!
  • 154275 09/09/05 09:54 Moffit New "empty" encoder position
  • 154274 09/09/05 09:45 Moffit Until Further Notice: No targets above BeO
  • 154273 09/09/05 09:30 Moffit Empty target position doesn't appear so empty..
  • 154272 09/09/05 09:25 Moffit BPM oscillators off
  • 154270 09/09/05 08:53 R. Michaels turn stuff on, checks
  • 154267 09/09/05 08:34 atrig Trigger-Download
  • 154265 09/09/05 01:01 Moffit Start of Parity Run 4955
  • 154264 09/09/05 00:59 moffit Run end 4954
  • Thursday

  • 154263 09/08/05 17:50 H. Benaoum Regression & Dithering Slopes for Detectors vs. Slug (1-76) without slugs 1,2,3,39,40 & 41.
  • 154262 09/08/05 17:39 R. Michaels Hall A swept
  • 154257 09/08/05 08:24 R. Michaels weekend schedule
  • Wednesday

  • 154252 09/07/05 20:32 lkaufman BPM oscillators on
  • 154248 09/07/05 19:57 lkaufman Helicity control signals working again
  • Tuesday

  • 154245 09/06/05 17:44 paschke helicity control signals not yet working
  • 154242 09/06/05 16:27 Moffit Re-analyzed standard rootfiles slug1-46
  • 154241 09/06/05 15:16 H. Benaoum Raw, Regressed, Dithering Corrected and Regression Corrected Detectors Normalized Asymmetries versus Slug Number (1-76) without slugs 39,40 and 41.
  • Saturday

  • 154238 09/03/05 00:35 Moffit BPM oscillators off
  • Friday

  • 154235 09/02/05 19:38 lkaufman WAC Daily Summary for Sept. 2
  • 154234 09/02/05 19:30 lkaufman WAC Summary for Slug 76
  • 154233 09/02/05 19:28 lkaufman WAC Summary for Slug 75
  • 154232 09/02/05 13:51 Moffit Detector LED electronics installed
  • 154228 09/02/05 12:52 R. Michaels Cavity results at low current
  • 154224 09/02/05 11:42 R. Michaels Coulomb meter: 60C
  • 154223 09/02/05 11:28 Moffit HAPPEX DAQ raw data to adaql1/data3
  • 154216 09/02/05 08:53 Moffit BPM calibration oscillators on
  • 154215 09/02/05 08:40 R. Michaels shutdown notes.
  • 154214 09/02/05 08:18 H. Benaoum Owl Summary Shift
  • 154210 09/02/05 07:35 H. Benaoum Raw, Regressed, Dithering Corrected and Regression Corrected Detectors Normalized Asymmetries versus Slug Number (1-74) without slugs 39,40 and 41.
  • 154209 09/02/05 07:20 Sarah Hall A Tools Screen
  • 154208 09/02/05 07:11 H. Benaoum Target Sreen Capture before Warm up
  • 154203 09/02/05 05:47 lkaufman Low current cavity results from HAPPEX DAQ
  • 154198 09/02/05 04:47 H. Benaoum Right Arm is Back
  • 154195 09/02/05 03:46 H. Benaoum Right Arm is Back
  • 154194 09/02/05 03:11 H. Benaoum NetWork Problem : lost both arm & target
  • 154188 09/02/05 01:00 Sarah IHWP is now OUT!
  • 154186 09/02/05 00:52 lkaufman WAC Daily Summary for Sept. 1
  • 154185 09/02/05 00:09 lkaufman Swing Shift Summary
  • 154184 09/02/05 00:08 lkaufman Profile Scan: Run 4928
  • Thursday

  • 154174 09/01/05 21:34 lkaufman Beamline HV Alarm
  • 154173 09/01/05 21:14 Kelleher Hall A Tools
  • 154172 09/01/05 21:08 Kelleher target screens
  • 154171 09/01/05 21:02 lkaufman WAC Summary for Slug 74
  • 154168 09/01/05 20:10 Kelleher IHWP is now IN
  • 154163 09/01/05 18:20 lkaufman Cavities are ok!
  • 154160 09/01/05 18:13 Moffit Linearity test, take 2
  • 154159 09/01/05 18:11 lkaufman ROC 26 rebooted itself
  • 154157 09/01/05 17:41 lkaufman Low I cavity commis. part 3
  • 154154 09/01/05 17:26 lkaufman spot++
  • 154152 09/01/05 17:25 lkaufman Cavities back to nominal
  • 154147 09/01/05 17:03 lkaufman cavity monitor parameters for run 2604/run4919
  • 154141 09/01/05 16:52 lkaufman Cavity Monitor parameters for 50 nA running
  • 154138 09/01/05 16:39 R. Michaels Low I cavity commis. part 2
  • 154135 09/01/05 16:04 Saha Sift Summary
  • 154133 09/01/05 15:53 r gilman target screen capture for shift checklist
  • 154132 09/01/05 15:53 Saha Hall A Tools Screen
  • 154121 09/01/05 15:25 lkaufman Raster off w/ 1uA beam on He
  • 154112 09/01/05 15:04 R. Michaels cavity low-I commiss. part 1
  • 154111 09/01/05 15:02 lkaufman WAC Summary for Slug 73
  • 154104 09/01/05 13:38 paschke Beam studies: helicity magnet test
  • 154103 09/01/05 13:31 R. Snyder 4A & 4B position differences through 72 slugs
  • 154095 09/01/05 12:55 R Gilman switched target to Helium 20 cm
  • 154093 09/01/05 12:33 paschke helicity signal to helicity magnets reconnected
  • 154090 09/01/05 12:28 paschke starting beam studies
  • 154086 09/01/05 10:11 R Gilman target change
  • 154081 09/01/05 08:40 Ole Hansen HWP to OUT
  • 154079 09/01/05 08:02 H. Benaoum Owl Shift Summary
  • 154073 09/01/05 06:38 H. Benaoum Spot++
  • 154071 09/01/05 06:25 H. Benaoum Optics Runs
  • 154070 09/01/05 05:50 H. Benaoum Moved to Helium
  • 154068 09/01/05 05:39 H. Benaoum Target Screen
  • 154067 09/01/05 05:38 Sarah Hall A Tools Snapshot
  • 154061 09/01/05 05:09 H. Benaoum Moved to Tantalum
  • 154060 09/01/05 05:09 H. Benaoum Moved to Carbon
  • 154051 09/01/05 04:32 H. Benaoum Compton Quarter Wave Plate Scan
  • 154050 09/01/05 04:31 H. Benaoum Moved to BeO
  • 154048 09/01/05 04:24 R. Feuerbach Direction of +X on BPM4B
  • 154044 09/01/05 04:11 H. Benaoum Moved to Helium
  • 154039 09/01/05 03:46 H. Benaoum Optics Target
  • 154024 09/01/05 02:45 H. Benaoum Moved to Helium Target
  • 154019 09/01/05 02:18 H. Benaoum Moved to Optics Target
  • 154014 09/01/05 01:56 H. Benaoum Moved to Carbon
  • 154013 09/01/05 01:55 H. Benaoum Errata : Compton Quarter Wave Plate Scan(08/31/05)
  • 154011 09/01/05 01:47 H. Benaoum Compton Quarter Wave Plate Scan (08/31/05)
  • 154005 09/01/05 01:13 H. Benaoum Moved to Tantalum
  • 154001 09/01/05 01:02 lkaufman WAC Daily Summary for Aug. 31
  • 153996 09/01/05 00:37 H. Benaoum Moved to BeO
  • 153992 09/01/05 00:01 lkaufman Swing Shift Summary